3/5 = 1(x - 1/3 - x/5)
Answer by algebrapro18(206)

well since since the right side is only being multiplied by 1 we don't need to distribute so you are left with
3/5 = x-1/3-x/5
now I really dislike fracitons so I clear them by multiplying by the GCF. This will get you to
9 = 15x-5-3x
now collect like terms
9 = 12x-5
now add 5 to both sides
14=12x
now divide both sides by 12
14/12=x
7/6 = x
checking you get
3/5 = 7/6 - 1/3 - (7/6)/5
3/5 = 5/6 - 7/30
3/5 = (25-7)/30
3/5 = 18/30
3/5 = 3/5

If you have any questions feel free to e-mail me at dettmanb@uwplatt.edu